# LAMBench: A Benchmark for Large Atomic Models

### v1

@@ Abstract
Large atomic models (LAMs) have undergone remarkable progress recently, emerging as universal or fundamental representations of the potential energy surface defined by the first-principles calculations of atomic systems. However, our understanding of the extent to which these models achieve true universality, as well as their comparative performance across different models, remains limited. This gap is largely due to the lack of comprehensive benchmarks capable of evaluating the effectiveness of LAMs as approximations to the universal potential energy surface. In this study, we introduce LAMBench, a benchmarking system designed to evaluate LAMs in terms of their generalizability, adaptability, and applicability. These attributes are crucial for deploying LAMs as ready-to-use tools across a diverse array of scientific discovery contexts. We benchmark eight state-of-the-art LAMs released prior to April 1, 2025, using LAMBench. Our findings reveal a significant gap between the current LAMs and the ideal universal potential energy surface. They also highlight the need for incorporating cross-domain training data, supporting multi-fidelity modeling, and ensuring the models’ conservativeness and differentiability. As a dynamic and extensible platform, LAMBench is intended to continuously evolve, thereby facilitating the development of robust and generalizable LAMs capable of significantly advancing scientific research. The LAMBench code is open-sourced at https://github.com/deepmodeling/lambench, and an interactive leaderboard is available at https://www.aissquare.com/openlam?tab=Benchmark.
@@ I Introduction
The widespread adoption of large language models (LLMs) is largely driven by the development of general-purpose foundation models pretrained on vast and diverse corpora covering a wide range of disciplines and topics[1]. These foundation models are feasible because there exist common patterns to learn — namely, the shared logic of human language — despite its apparent diversity. In the field of molecular modeling, the fundamental physical principles of quantum mechanics, particularly the Schrödinger equation[2], apply universally to all atomic systems, assuming that relativistic effects are negligible. Under the Born-Oppenheimer approximation[3] a universal potential energy surface (PES) is defined as the ground state solution of the electronic Schrödinger equation, with the nuclear positions treated as input parameters. Consequently, it is feasible to develop a foundational machine learning model to approximate this universal PES. We refer to these molecular foundation models as Large Atomic Models (LAMs) to emphasize their role in capturing fundamental atomic and molecular interactions across diverse chemical systems[4]. LAMs are typically developed through a two-stage process: an initial pretraining phase on broad, diverse atomic datasets to learn a latent representation of the universal PES, followed by fine-tuning on specific downstream datasets to specialize the model for particular target applications.
@@
Nevertheless, domain-specific LAMs are advancing rapidly. For example, MACE-MP-0[11] and SevenNet-0[12] are both trained on the MPtrj dataset[13] from the Inorganic Materials domain at the PBE/PBE+U level of theory. AIMNet[14] and Nutmeg[15] are trained in the domain of small molecules at the SMD(Water)-$\omega$B97X/def2-TZVPP and the $\omega$B97M- D3(BJ)/def2-TZVPPD level of theory, respectively. The rapid advancement of these domain-specific LAMs has transformed the field of atomistic modeling, offering powerful tools for understanding complex inorganic materials and bio-molecular systems. To fully harness the diverse training data from various research domains and maximize the potential of LAMs in learning universal PES, the multitask pretraining strategy presents a promising approach. This strategy encodes shared knowledge into a unified structure with high representational capacity while integrating domain-specific components into multiple neural networks with relatively lower representational power[16, 4]. However, determining the extent to which multitask-trained LAMs approach a truly universal PES remains a challenging question.
@@
In the field of molecular modeling, existing benchmarks exhibit two significant limitations. Firstly, they are intrinsically domain-specific, concentrating on isolated sub-fields rather than encompassing a variety of atomic systems. For instance, datasets such as QM9[22] and MD17[23] are used to benchmark molecular property predictions and molecular dynamics (MD) trajectories of small molecules, respectively. These benchmarks are predominantly employed to assess machine learning models within chemical science. The Matbench Discovery[24] evaluates models in the Inorganic Materials domain based on their ability to predict material stability. The Open Catalyst challenges[25] assess models on predicting adsorption energies and relaxed structures for various adsorbate-catalyst combinations. While these benchmarks have played a crucial role in advancing domain-specific LAMs, their fragmented approach undermines the pursuit towards the universal PES model. Secondly, existing assessment methods often fail to reflect real-world application scenarios, reducing their relevance to scientific discovery and technological innovation. For instance, conventional evaluation metrics based on static test sets may not adequately capture the true performance of a model in tasks requiring physically meaningful energy landscapes[26]. Specifically, non-conservative models - where atomic forces are directly inferred from neural networks rather than obtained from the gradient of the predicted energy[27] — can exhibit high apparent accuracy but struggle in applications demanding strict energy conservation, such as MD simulations[28]. The MLIP-Arena benchmark[29] is a step in the direction toward bridging this gap, emphasizing the practical usability of LAMs in tasks such as MD stability and physical property predictions. However, it places less emphasis on evaluating a model’s capacity to generalize across diverse atomic systems or adapt to tasks beyond its training scope, both of which are essential for assessing the performance of LAMs in real scientific discovery.
@@
To address these limitations, we introduce LAMBench, a comprehensive benchmark system designed to rigorously evaluate LAMs across domains, simulation regimes, and application scenarios. Employing LAMBench, we evaluated the performance of eight prominent LAMs released prior to April 1, 2025, uncovering a significant discrepancy between these models and the universal PES. Our findings suggest that enhancing LAM performance requires simultaneous training with data from a diverse array of research domains. Additionally, supporting multi-fidelity at inference time is essential to satisfy the varying requirements of exchange-correlation functionals across different domains. It is also critical to maintain the model’s conservativeness and differentiability to optimize performance in property prediction tasks and ensure stability in molecular dynamics simulations. We believe that the introduction of LAMBench will significantly expedite the development of LAMs, facilitating the creation of ready-to-use models that enhance the pace of real scientific discovery.
@@ II.1 The LAMBench system
The LAMBench system is designed to benchmark diverse Large Atomic Models (LAMs) across multiple tasks within a high-throughput workflow, with automation integral to task calculation, result aggregation, analysis, and visualization, as depicted in Figure 1. The implementation details of LAMBench are elaborated in Section IV.2. Key to the system’s effectiveness are the design of the benchmark tasks and the methodologies employed for result interpretation. These benchmark tasks are developed to assess three fundamental capabilities of an LAM: generalizability, adaptability, and applicability. Generalizability pertains to the accuracy of an LAM when utilized as a universal potential across a diverse range of atomic systems. Adaptability denotes the LAM’s capacity to be fine-tuned for tasks beyond potential energy prediction, with particular emphasis on structure-property relationship tasks. Applicability, on the other hand, concerns the stability and efficiency of deploying LAMs in real-world simulations.
@@
It is important to note that there remains no consensus on the precise definition of OOD or the criteria by which two distributions are considered different. Some researchers define OOD data as those exploring different configurational spaces[30, 31], while others highlight differences in chemical space as critical[25, 32]. In this study, we adopt a practical approach by considering OOD test datasets as downstream datasets designed to address specific scientific challenges, such as training a machine learning potential model for simulating carbon deposition on metal surfaces[33]. These scenarios are most compatible with the downstream applications of LAMs.
@@
In LAMBench, the generalizability is quantitatively assessed using two types of tasks, the force field task and the property calculation task. The force field task assesses the accuracy of LAMs in predicting energy, force, and the virial tensor (when periodic boundary conditions are applied) using 17 datasets that have been utilized in the literature to train machine learning potentials for addressing scientific challenges. Detailed information on these datasets is provided in LABEL:tab:datasets. These test datasets encompass five distinct scientific domains: Inorganic Materials, Catalysis, Reactions, Small Molecules, and Biomolecules and Supramolecules. Given that energy labels calculated via DFT can vary by an arbitrary constant due to variations in pseudopotential selection and software implementations, LAMs are consistently used to predict the energy difference between the label and a dummy model that estimates potential energy solely based on the chemical formula. In contrast, force and virial predictions are directly obtained from the LAMs, as the dummy model invariably yields null predictions for force and virial.
@@
To evaluate the generalizability of LAMs, it is crucial to establish a comprehensive performance indicator that integrates all the energy, force, and virial errors calculated for each configuration across every dataset and domain. An arithmetic mean of the error data is inappropriate due to the differing units of the quantities involved. To address this challenge, we introduce a dimensionless error metric, $\bar{M}^{m}_{\mathrm{FF}}$, with detailed information available in Section IV.3. The error metric is structured so that a dummy model yields a value of $\bar{M}^{m}_{\mathrm{FF}}=1$, whereas an ideal model that perfectly aligns with DFT labels achieves a metric of $\bar{M}^{m}_{\mathrm{FF}}=0$. A lower value of this error metric signifies enhanced generalizability.
@@
Achieving high accuracy in the force field task does not necessarily ensure enhanced performance in property calculations when utilizing the LAM as the force field[26]. For instance, predicting the bulk and shear moduli necessitates finite difference approximations of the second-order derivatives of the potential energy. Consequently, models lacking smoothness up to the second-order derivatives may not consistently yield accurate predictions, even if they perform well in force field tasks. Given the significant variation in investigated properties across different research domains, LAMBench offers a flexible and modular design (see Section IV.2). This design facilitates the implementation of property calculation benchmark tasks, enabling them to be tailored to the specific requirements of diverse research fields.
@@
In this study, we have adapted the MDR Phonon benchmark[34] to evaluate the performance of LAM models in computing key phonon properties, including maximum phonon frequency ($\omega_{\max}$), entropy ($S$), free energy ($F$), and heat capacity at constant volume ($C_{V}$). Additionally, we have utilized the TorsionNet500 benchmark to assess the capability of LAMs in calculating the torsion energy profile of typical drug-like fragments[35]. The Phonon and TorsionNet500 benchmarks are classified as properties of interest within the Inorganic Materials and Small Molecules domains, respectively. These adaptations serve as prototypes for evaluating downstream generalizability on domain-specific property calculations. Further tasks targeting Reactions, Catalysis and other research domains are under active development to align with the evolving capabilities of LAMs. Similar to the approach taken in the force field task, we have introduced a dimensionless error metric, denoted as $\bar{M}^{m}_{\mathrm{PC}}$. This metric is designed to assess the generalizability of an LAM in the property calculation task. For comprehensive details, please refer to Section IV.3.
@@
Quantifying the inference efficiency of LAMs in a rigorous and meaningful manner requires careful design, as the observed efficiency can be highly structure-dependent, and different LAMs may exhibit varying performance on the same input. LAMs are designed to be applicable to tasks involving systems of varying sizes, from evaluating the conformational energy of molecules composed of dozens of atoms to large-scale simulations of viruses involving millions of atoms[38]. The efficiency of LAMs—measured as the computational time consumed per atom while determining the energy, forces, and virial of atomic systems—is significantly influenced by the system size and the extent to which the many-core parallelism of modern GPUs is utilized. For relatively small systems, atom-wise efficiency is often reduced due to limited opportunities for parallelization, even though they require less overall computational time for evaluation. As system size increases to approximately 1000 atoms, the average inference time per atom tends to stabilize, as illustrated by Supporting Information Figure S-3. For substantially larger systems, exceeding the memory capacity of a single GPU, multi-GPU parallelism becomes essential. However, such systems are not ideal benchmarks for assessing the efficiency of LAMs, as they are generally computationally intensive, and multi-GPU parallelism is not supported by all the LAMs.
@@
To address the significant size-dependency issues when measuring efficiency in small systems, we concentrate on atomic systems that are large enough to achieve converged efficiency yet small enough to be computed without the need for multi-GPU parallelism. To this end, we randomly sample 1,600 structures from the Inorganic Materials and Catalysis domains in the aforementioned force field task test sets. Each structure is subjected to periodic boundary conditions and duplicated to contain between 800 and 1,000 atoms. This range is sufficiently large to approach the convergence regime while remaining within the memory constraints of most models, thereby avoiding out-of-memory (OOM) errors. Other domains, such as Small Molecules, are excluded from this evaluation, as non-periodic systems generally do not pose sufficient computational demands to effectively stress the inference capabilities of LAMs.
@@
To facilitate the comparison of efficiency across models, we propose a dimensionless efficiency metric, denoted as $M_{\mathrm{E}}^{m}$. This metric is defined as the normalized inverse of the averaged inference time measured for the 1,600 structures. The detailed definition is provided in Section IV.4. A value of $M_{\mathrm{E}}^{m}=1$ corresponds to an efficiency equivalent to a reference value of 0.01 $(\mathrm{\mu s/atom})^{-1}$, with higher value indicating greater efficiency. It is crucial to note that the calculator is fully warmed up by processing 400 structures prior to recording the inference time for the 1,600 structures.
@@
We assessed the stability of LAMs by examining energy conservation in NVE ensemble (microcanonical ensemble) MD simulations across nine atomic systems. These systems were randomly selected from diverse domains, including four periodic structures from the Materials Project[39], three molecular systems from the SPICE2 dataset[40], and two catalytic surface systems from the OC2M dataset[41], as detailed in Supporting Information Figure S-1. For each system, a 10 ps NVE-MD simulation was conducted with a timestep of 1 fs. The initial atomic velocities were sampled from the Boltzmann ensemble at 300 K. The drift of total energy along MD trajectories was quantified using the slope derived from a linear regression applied to the total energy history, with the initial 2 ps of simulation excluded as a warm-up period. Stability is assessed via the instability metric, $M^{m}_{\mathrm{IS}}$, which is calculated based on total energy drift, where smaller values denote improved stability. A detailed definition is provided in Section IV.4.
@@
Table 1: Summary of datasets utilized in force field generalizability assessments. The table details the domain of each test dataset, available labels (E for energy, F for force, and V for virial), number of configurations (frames), exchange-correlation (XC) functional applied for labeling, and additional descriptions. Details on data cleaning procedures can be found in Supporting Information Section S-3.
Name Domain Labels Frames Level of XC Description
@@
Lopanitsyna2023Modeling_A[43] Inorganic Materials EFV 450 PBE-sol Commonly known as HEA25S. A dataset of high-entropy alloy surfaces, focusing on d-block transition metals covering 25 elements.
@@
Lopanitsyna2023Modeling_B[44] Inorganic Materials EFV 25628 PBE-sol Commonly known as HEA25. A dataset of high-entropy alloy bulk structures, focusing on d-block transition metals covering 25 elements.
Dai2024Deep[45] Inorganic Materials EFV 2842 PBE A dataset of high-entropy transition metal diboride (HEMB2) and carbide (HEMC) ceramics.
@@
Torres2019Analysis[46] Inorganic Materials EF 4378 PBE A dataset of Ca-bearing minerals, focusing on silicates and carbonates.
@@
Sours2023Applications[47] Inorganic Materials EF 17414 PBE-D3(BJ) Dataset consisting of 219 different pure silica zeolite topologies.
@@
Batzner2022equivariant[48] Inorganic Materials EF 7499 PBE Down-sampled dataset containing lithium phosphate amorphous glass used by NequIP graph neural network models as evaluation configurations.
WBM_25k[49] Inorganic Materials E 25696 PBE A subset of the WBM dataset randomly down-sampled to 25,696 frames.
Subalex_9k[50] Inorganic Materials EFV 9157 PBE A subset of sAlexandria dataset, randomly down-sampled to 9,157 frames.
@@
ANI-1x[51] Small Molecules EF 8861 $\omega$B97X / 6-31G* A down-sampled dataset from the training data of the ANI-1x potential, containing organic molecule structures.
Zhang2024Active[33] Catalysis EF 452 PBE-D3 Dynamic simulations of carbon deposition on metal surfaces.
Zhang2019Bridging[52] Catalysis EF 9842 PBE Interaction of carbon dioxide with a movable Ni(100) surface.
Vandermause2022Active[53] Catalysis EFV 250 PBE Direct simulation of hydrogen turnover on Pt(111) catalyst surfaces.
Villanueva2024Water[54] Catalysis EF 14859 PBE-D3 Selective CO2 hydrogenation to methanol over oxide catalysts.
Guan2022Benchmark[55] Reactions EF 17412 $\omega$B97X-V / cc-pVTZ Dataset of hydrogen combustion reactions covering 19 reaction channels.
Gasteiger2020Fast[56] Reactions EF 9480 revPBE-D3 / def2-TZVP Validation set from molecular collision experiments with small organic molecules.
MD22[57] Biomolecules EF 2122 PBE+MBD / tighta Molecular dynamics (MD) trajectories of 42-atom tetrapeptide Ac-Ala3-NHMe.
@@
AIMD-Chig[58] Biomolecules EF 19800 M06-2X / 6–31G* A down-sampled MD dataset containing conformations of chignolin protein.
a The “tight” setting corresponds to the default high-accuracy mode in FHI-aims.[42]

### v2

# LAMBench: A Benchmark for Large Atomistic Models

@@ Abstract
Large Atomistic Models (LAMs) have undergone remarkable progress recently, emerging as universal or fundamental representations of the potential energy surface defined by the first-principles calculations of atomistic systems. However, our understanding of the extent to which these models achieve true universality, as well as their comparative performance across different models, remains limited. This gap is largely due to the lack of comprehensive benchmarks capable of evaluating the effectiveness of LAMs as approximations to the universal potential energy surface. In this study, we introduce LAMBench, a benchmarking system designed to evaluate LAMs in terms of their generalizability, adaptability, and applicability. These attributes are crucial for deploying LAMs as ready-to-use tools across a diverse array of scientific discovery contexts. We benchmark ten state-of-the-art LAMs released prior to August 1, 2025, using LAMBench. Our findings reveal a significant gap between the current LAMs and the ideal universal potential energy surface. They also highlight the need for incorporating cross-domain training data, supporting multi-fidelity modeling, and ensuring the models’ conservativeness and differentiability. As a dynamic and extensible platform, LAMBench is intended to continuously evolve, thereby facilitating the development of robust and generalizable LAMs capable of significantly advancing scientific research. The LAMBench code is open-sourced at https://github.com/deepmodeling/lambench, and an interactive leaderboard is available at https://www.aissquare.com/openlam?tab=Benchmark.
@@ I Introduction
The widespread adoption of large language models (LLMs) is largely driven by the development of general-purpose foundation models pretrained on vast and diverse corpora covering a wide range of disciplines and topics[1]. These foundation models are feasible because there exist common patterns to learn — namely, the shared logic of human language — despite its apparent diversity. In the field of molecular modeling, the fundamental physical principles of quantum mechanics, particularly the Schrödinger equation[2], apply universally to all atomistic systems, assuming that relativistic effects are negligible. Under the Born-Oppenheimer approximation[3] a universal potential energy surface (PES) is defined as the ground state solution of the electronic Schrödinger equation, with the nuclear positions treated as input parameters. Consequently, it is feasible to develop a foundational machine learning model to approximate this universal PES. We refer to these molecular foundation models as Large Atomistic Models (LAMs) to emphasize their role in capturing fundamental atomic and molecular interactions across diverse chemical systems[4]. LAMs are typically developed through a two-stage process: an initial pretraining phase on broad, diverse atomic datasets to learn a latent representation of the universal PES, followed by fine-tuning on specific downstream datasets to specialize the model for particular target applications.
@@
[…]
@@
In the field of molecular modeling, existing benchmarks exhibit two significant limitations. Firstly, they are intrinsically domain-specific, concentrating on isolated sub-fields rather than encompassing a variety of atomistic systems. For instance, datasets such as QM9[23] and MD17[24] are used to benchmark molecular property predictions and molecular dynamics (MD) trajectories of small molecules, respectively. These benchmarks are predominantly employed to assess machine learning models within chemical science. The Matbench Discovery[25] evaluates models in the Inorganic Materials domain based on their ability to predict material stability. The Open Catalyst challenges[26] assess models on predicting adsorption energies and relaxed structures for various adsorbate-catalyst combinations. While these benchmarks have played a crucial role in advancing domain-specific LAMs, their fragmented approach undermines the pursuit towards the universal PES model. Secondly, existing assessment methods often fail to reflect real-world application scenarios, reducing their relevance to scientific discovery and technological innovation. For instance, conventional evaluation metrics based on static test sets may not adequately capture the true performance of a model in tasks requiring physically meaningful energy landscapes[27]. Specifically, non-conservative models – where atomic forces are directly inferred from neural networks rather than obtained from the gradient of the predicted energy[28] – can exhibit high apparent accuracy but struggle in applications demanding strict energy conservation, such as MD simulations[29]. The MLIP-Arena benchmark[30] is a step in the direction toward bridging this gap, emphasizing the practical usability of LAMs in tasks such as MD stability and physical property predictions. However, it places less emphasis on evaluating a model’s capacity to generalize across diverse atomistic systems or adapt to tasks beyond its training scope, both of which are essential for assessing the performance of LAMs in real scientific discovery.
@@
To address these limitations, we introduce LAMBench, a comprehensive benchmark system designed to rigorously evaluate LAMs across domains, simulation regimes, and application scenarios. Utilizing LAMBench, we assessed the performance of ten leading LAMs released before August 1, 2025, excluding the UMA models[17] due to licensing restrictions[31]. Our analysis revealed a substantial discrepancy between these models and the universal PES. Our findings suggest that enhancing LAM performance requires simultaneous training with data from a diverse array of research domains. Additionally, supporting multi-fidelity at inference time is essential to satisfy the varying requirements of exchange-correlation functionals across different domains. It is also critical to maintain the model’s conservativeness and differentiability to optimize performance in property prediction tasks and ensure stability in molecular dynamics simulations. We believe that the introduction of LAMBench will significantly expedite the development of LAMs, facilitating the creation of ready-to-use models that enhance the pace of real scientific discovery.
@@ II.1 The LAMBench system
The LAMBench system is designed to benchmark diverse Large Atomistic Models (LAMs) across multiple tasks within a high-throughput workflow, with automation integral to task calculation, result aggregation, analysis, and visualization, as depicted in Figure 1. The implementation details of LAMBench are elaborated in Section IV.3. Key to the system’s effectiveness are the design of the benchmark tasks and the methodologies employed for result interpretation. These benchmark tasks are developed to assess three fundamental capabilities of an LAM: generalizability, adaptability, and applicability. Generalizability pertains to the accuracy of an LAM when utilized as a universal potential across a diverse range of atomistic systems. Adaptability denotes the LAM’s capacity to be fine-tuned for tasks beyond potential energy prediction, with particular emphasis on structure-property relationship tasks. Applicability, on the other hand, concerns the stability and efficiency of deploying LAMs in real-world simulations. In this study, all tasks are defined, and all results are obtained using LAMBench version v0.3.1.
@@
It is important to note that there remains no consensus on the precise definition of OOD or the criteria by which two distributions are considered different. Some researchers define OOD data as those exploring different configurational spaces[32, 33], while others highlight differences in chemical space as critical[26, 34]. In this study, we adopt a practical approach by considering OOD test datasets as downstream datasets designed to address specific scientific challenges[35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. These scenarios are most compatible with the downstream applications of LAMs.
@@
In LAMBench, the generalizability is quantitatively assessed using two types of tasks, the force field task and the property calculation task. The force field task measures the accuracy of LAMs in predicting energy, force, and the virial tensor (under periodic boundary conditions) using twelve downstream datasets employed in the literature to train machine learning potentials for solving scientific problem, without any fine-tuning. Detailed information on these datasets is provided in LABEL:tab:datasets. These test datasets encompass three distinct scientific domains: Inorganic Materials, Catalysis, and Molecules.
It is important to note that the majority of LAMs are trained on datasets labeled with PBE/PBE+U exchange-correlation (XC) functionals, which may not align with the XC functional used in the test datasets. To address this mismatch when evaluating the generalizability of LAMs, we relabeled five datasets at the PBE level, as detailed in the Methods section Section IV. Dispersion corrections were added through post-prediction process with the DFTD3-ASE interface[47, 48], when necessary. Minor mismatches in pseudopotentials, basis sets, and software implementations within DFT, which are less significant than the XC functional discrepancies, were not addressed in the current benchmark. Given that energy labels calculated via DFT can vary by an arbitrary constant, LAMs are consistently used to predict the energy difference between the label and a dummy model that estimates potential energy solely based on the chemical formula. In contrast, force and virial predictions are directly obtained from the LAMs, as the dummy model invariably yields null predictions for force and virial.
@@
Establishing a comprehensive performance indicator that synthesizes energy, force, and virial errors for each configuration across all datasets and domains presents a complex challenge. An arithmetic mean of the error data is inappropriate due to the differing units of the quantities involved. To address this issue, we introduce a dimensionless error metric, $\bar{M}^{m}_{\mathrm{FF}}$, which essentially averages the relative errors in predictions across all datasets. Detailed information on this metric can be found in Section IV.4. The metric is structured so that a dummy model yields a value of $\bar{M}^{m}_{\mathrm{FF}}=1$, whereas an ideal model that perfectly aligns with DFT labels achieves a metric of $\bar{M}^{m}_{\mathrm{FF}}=0$. A lower value of this error metric signifies enhanced generalizability.
In the error metric, the relative error is preferred over the absolute error to better reflect a model’s accuracy in relation to the task’s difficulty. For example, the standard deviations in energy labels for the Lopanitsyna2023Modeling and MD22 datasets are 511 meV/atom and 9 meV/atom, respectively. This considerable difference in scales indicates the greater diversity of configurations in the Lopanitsyna2023Modeling dataset. If a model exhibits the same energy RMSE of 9 meV/atom for both datasets, it achieves a 2 % error in the Lopanitsyna2023Modeling dataset, while demonstrating no accuracy in the MD22 dataset. Therefore, given the substantial differences between datasets, the relative error offers a more meaningful measure of model performance.
@@
Achieving high accuracy in the force field task does not necessarily ensure enhanced performance in property calculations when utilizing the LAM as the force field[27]. For instance, predicting the bulk and shear moduli necessitates finite difference approximations of the second-order derivatives of the potential energy. Consequently, models lacking smoothness up to the second-order derivatives may not consistently yield accurate predictions, even if they perform well in force field tasks. Given the significant variation in investigated properties across different research domains, LAMBench offers a flexible and modular design (see Section IV.3). This design facilitates the implementation of property calculation benchmark tasks, enabling them to be tailored to the specific requirements of diverse research fields.
@@
In this study, we include property calculation tasks across the aforementioned research domains. For the Inorganic Materials domain, we adapt the MDR Phonon benchmark [49] to assess the performance of LAM models in computing essential phonon properties, including maximum phonon frequency ($\omega_{\max}$), entropy ($S$), free energy ($F$), and heat capacity at constant volume ($C_{V}$). Additionally, we employ the elasticity benchmark [50, 51] to evaluate the accuracy of LAM models in calculating the shear modulus (GVHR) and bulk modulus (KVHR). For the Molecules domain, we use the TorsionNet500 benchmark[52] to evaluate LAM performance in calculating torsion energy profiles of typical drug-like fragments, and the Wiggle150 benchmark[53] to assess energy profiles of highly strained conformers. For the Catalysis domain, we employ the OC20NEB-OOD dataset[54] to benchmark the ability of LAMs to predict energy barriers for three reaction types—transfer, dissociation, and desorption—using DFT-optimized structures along the NEB trajectories. For benchmarks in the Inorganic Materials domain, the PBE/PBE+U XC functional was adopted to compute the reference predictions. For the Molecules and Catalysis domains, reference data are obtained at the CCSD(T) and RPBE levels, respectively, which are considered to be more accurate than the PBE XC functional within their respective domains. The incorporation of these property calculation tasks serves as a prototype for assessing downstream generalizability in domain-specific property calculations. Additional tasks can be seamlessly integrated to align with the evolving capabilities of LAMs. In line with the approach used in the force field task, we introduce a dimensionless error metric, denoted as $\bar{M}^{m}_{\mathrm{PC}}$, to evaluate the generalizability of an LAM in property calculation tasks. For comprehensive details, please refer to Section IV.4.
@@
Quantifying the inference efficiency of LAMs in a rigorous and meaningful manner requires careful design, as the observed efficiency can be highly structure-dependent, and different LAMs may exhibit varying performance on the same input. LAMs are designed to be applicable to tasks involving systems of varying sizes, from evaluating the conformational energy of molecules composed of dozens of atoms to large-scale simulations of viruses involving millions of atoms[57]. The efficiency of LAMs—measured as the computational time consumed per atom while determining the energy, forces, and virial of atomistic systems—is significantly influenced by the system size and the extent to which the many-core parallelism of modern GPUs is utilized. For relatively small systems, atom-wise efficiency is often reduced due to limited opportunities for parallelization, even though they require less overall computational time for evaluation. As system size increases to approximately 1000 atoms, the average inference time per atom tends to stabilize, as illustrated by Supporting Information Figure S-3. For substantially larger systems, exceeding the memory capacity of a single GPU, multi-GPU parallelism becomes essential. However, such systems are not ideal benchmarks for assessing the efficiency of LAMs, as they are generally computationally intensive, and multi-GPU parallelism is not supported by all the LAMs.
@@
To address the significant size-dependency issues when measuring efficiency in small systems, we concentrate on atomistic systems that are large enough to achieve converged efficiency yet small enough to be computed without the need for multi-GPU parallelism. To this end, we randomly sample 1,000 structures from the Inorganic Materials and Catalysis domains in the aforementioned force field task test sets. Each structure is subjected to periodic boundary conditions and is replicated by a multiplier, dynamically determined using a binary search algorithm. This approach, with an upper limit of 1,000 atoms for efficiency, ensures that measurements are conducted in or near the convergence regime. The Molecules domain is excluded from this evaluation, as its systems contain only a few dozen atoms and therefore do not impose sufficient computational demands to meaningfully stress the inference capabilities of LAMs.
@@
To facilitate the comparison of efficiency across models, we propose a dimensionless efficiency metric, denoted as $M_{\mathrm{E}}^{m}$. This metric is defined as the normalized inverse of the average inference time measured for 900 out of the 1,000 randomly sampled structures. A detailed definition is provided in Section IV.5. A value of $M_{\mathrm{E}}^{m}=1$ corresponds to an efficiency equivalent to a reference value of 0.01 $(\mathrm{\mu s/atom})^{-1}$, with higher value indicating greater efficiency. It is important to note that the calculator is fully warmed up by processing 100 structures, which are excluded from the 1,000 randomly sampled structures, before recording the inference time on the remaining 900 structures.
@@
We assessed the stability of LAMs by examining energy conservation in NVE ensemble (microcanonical ensemble) MD simulations across nine atomistic systems. The extent of drift observed over time serves as a practical indicator of a model’s stability, specifically in maintaining a finite system energy in extended MD trajectories. These test systems were randomly selected from diverse domains, including four periodic structures from the Materials Project[58], three molecular systems from the SPICE2 dataset[59], and two catalytic surface systems from the OC2M dataset[60], as detailed in Supporting Information Figure S-1. For each system, a 10 ps NVE-MD simulation was conducted with a timestep of 1 fs. The initial atomic velocities were sampled from the Boltzmann ensemble at 300 K. The drift of total energy along MD trajectories was quantified using the slope derived from a linear regression applied to the total energy history, with the initial 2 ps of simulation excluded as a warm-up period. Stability is assessed via the instability metric, $M^{m}_{\mathrm{IS}}$, which is calculated based on total energy drift, where smaller values denote improved stability. A detailed definition is provided in Section IV.5.
@@
Table 1: Summary of datasets utilized in force field generalizability assessments. The table details the domain of each test dataset, available labels (E for energy, F for force, and V for virial), number of configurations (frames), exchange-correlation (XC) functional applied for labeling, and additional descriptions.
Name Domain Labels Frames Description
@@
Mazitov2024Surface[35] Inorganic Materials EFV 422 A down-sampled dataset of HEA25S, for high-entropy alloy surfaces, focusing on d-block transition metals covering 25 elements.
@@
Lopanitsyna2023Modeling[36] Inorganic Materials EFV 900 A down-sampled dataset of HEA25, for high-entropy alloy bulk structures, focusing on d-block transition metals covering 25 elements.
@@
Torres2019Analysis[37] Inorganic Materials EF 1000 A down-sampled dataset of Ca-bearing minerals, focusing on silicates and carbonates.
@@
Sours2023Predicting[38] Inorganic Materials EF 1000 A down-sampled Dataset consisting of 219 different pure silica zeolite topologies.
@@
Batzner2022equivariant[39] Inorganic Materials EF 1000 A down-sampled dataset containing lithium phosphate amorphous glass used by NequIP graph neural network models as evaluation configurations.
Gao2025Spontaneous[40] Inorganic Materials EFV 1000 A down-sampled Bilayer 2D MoS2 structures.
@@
ANI-1x[41] Molecules EF 997 A down-sampled dataset from the training data of the ANI-1x potential, containing organic molecule structures.
MD22[42] Molecules EF 1000 A down-sampled dataset from MD22, where molecular dynamics (MD) trajectories of 42-atom tetrapeptide Ac-Ala3-NHMe, docosahexaenoic acid, Stachyose, DNA base pair (AT-AT), DNA base pair (AT-AT-CG-CG).
@@
AIMD-Chig[43] Molecules EF 942 A down-sampled MD dataset containing conformations of chignolin protein.
Zhang2019Bridging[44] Catalysis EF 1000 A down-sampled dataset for interaction of carbon dioxide with a movable Ni(100) surface.
Vandermause2022Active[45] Catalysis EFV 250 A down-sampled dataset for direct simulation of hydrogen turnover on Pt(111) catalyst surfaces.
Villanueva2024Water[46] Catalysis EF 1000 A down-sampled dataset for selective CO2 hydrogenation to methanol over oxide catalysts.